1.) The vector A is given by the magnitude |A4.4 cm and the direction with respect to the +1-axis θΑ-140°. The vector B is given by the r-component Bz-4.4 cm and the y-component By3.2 cm                         b) Calculate the components of the vector C

c) Find the magnitude and direction of the vector C
Answer #1

1)

Ax = x-component of \underset{A}{\rightarrow} = |\underset{A}{\rightarrow}| Cos\theta _{A} = 4.4 Cos140 = - 3.37 cm

Ay = y-component of \underset{A}{\rightarrow} = |\underset{A}{\rightarrow}| Sin\theta _{A} = 4.4 Sin140 = 2.83 cm

Bx = 4.4 cm

By = 3.2 cm

Cx = Ax + Bx = - 3.37 + 4.4 = 1.03 cm

Cy = Ay + By = 2.83 + 3.2 = 6.03 cm

c)

using Pythagorean theorem

|C| = sqrt(Cx2 + Cy2) = sqrt((1.03)2 + (6.03)2) = 6.12 cm

direction : \theta = tan-1(Cy/Cx) = tan-1(6.03/1.03) = 80.31 deg